5.034 = 5.128*(1-Tax Rate)
How do I solve for the tax rate??

Answer by texttutoring(324)

I'm going to let Tax Rate = T for simplicity
5.034 = 5.128*(1-T)

Distribute:
5.034 = 5.128 - 5.128*T

Subtract 5.128 from both sides:
-0.094 = -5.128*T

Divide both sides by -5.128

T= 0.01833

If you want the rate in percent, multiply it by 100.

So the tax rate is 1.83%.